### v1

# The precision measurement of the electron anti-neutrino spectrum in beta-decay of nuclei

Abstract
The 144Ce – 144Pr electron antineutrino source is one of the most suitable radiochemical sources for experiments searching for the oscillations of active neutrinos to the light sterile state. In the current work the -spectra of – source have been measured by two types of -spectrometers based on silicon Si(Li) detectors in order to determine the energy spectrum of electron anti-neutrino emitted in the -decay of nuclei. The nuclear form factor of the ground state beta transition in has been obtained with high precision: . The reduced cross section for the inverse beta decay reaction on hydrogen for the electron anti-neutrino source has been defined as that provides sufficient sensitivity to search for a sterile neutrino with a mass of eV and a mixing angle using standard disappearance method.
I INTRODUCTION
The three-flavor neutrino () oscillations have been reliably observed using different neutrino sources and detectors. This fundamental discovery has demonstrated the presence of nonzero neutrino masses and has already went beyond the Standard Model (SM). Three mixing angles () and two mass difference squares of the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix have already been measured with acceptable accuracy [1]. However, the possible existence of additional sterile neutrino that mixes with the three SM neutrinos has remained one of the crucial unresolved problems of particle physics since the beginning of the century. The results of several neutrino experiments can be interpreted as indications of the light sterile neutrinos existence with of around [2].
The LSND experiment was the first one to observe excess of electron anti-neutrinos in the beam of neutrinos produced by pions decay at rest [3, 4]. The LSND results were tested by the MiniBooNE experiment which has also reported a low energy excess in and channels in neutrino beam from decay-in-flight pions. The overall significance of the MiniBooNE excess in both neutrino and antineutrino modes is [5, 6, 7]. The positive results of accelerator-based neutrino experiments LSND and MiniBooNE are currently being tested by JSNS2 [8] and MicroBooNE [9, 10] collaborations.
The “reactor antyneutrino anomaly” (RAA) has shown a discrepancy between the expected (based on a new anti-neutrino flux calculations [11, 12]) and the observed count rates of inverse beta-decay (IBD) reaction at the detectors installed close, m, to the nuclear reactors. This effect has been explained via mixing between electron and sterile (anti)neutrino with oscillation parameters and [13, 14]. However, the recent results of reactor experiments [15, 16, 17], as well as the new data on the spectrum of reactor neutrinos [18] probably do not confirm the existence of the RAA and its oscillation solution [19].
The next indication of the possible existence of a sterile neutrino has been related with () lack of count rate in Ga-Ge radiochemical solar neutrino detectors SAGE [20, 21] and GALLEX/GNO [22, 23] during calibration with intense artificial neutrino sources 51Cr and 37Ar. Recently, BEST collaboration has presented the new results obtained with source and Ga-Ge detector and the deficit of the observed reaction events stands at more than level [24, 25].
Several experiments with electron antineutrino sources (, , and ) were proposed to verify the detected anomalies by looking for neutrino oscillations at short distances [26, 27, 29]. One of these experiments was Borexino-SOX (Short distance neutrino Oscillations with BoreXino) in the Gran Sasso underground laboratory (Italy) with electron antineutrino () source [27, 28, 30, 31, 32]. The unprecedentedly low radioactive background of the Borexino detector, good energy and spacial resolutions together with large size allowing to perform an oscillometry (wave) measurement within the detector volume, and the well demonstrated capability to detect low energy electron antineutrinos have represented an ideal case for the study of short distance neutrino oscillations with artificial sources. Unfortunately, it was decided to suspend the Borexino_SOX experiment, however, the source remains one of the most promising for the future experiments searching for light sterile neutrinos.
The expected sensitivity of the experiment with source relative to the sterile neutrino oscillation parameters depends on the accuracy of flux determination as well as the energy spectrum shape for the electron anti-neutrino emitted in the decay of nuclei. The purpose of this work was a precise measurement of the spectra from source and determination of the spectrum of electron anti-neutrinos emitted in the decay nuclei with great precision.
II EXPERIMENTAL SETUP
The spectra of the source were measured by two setups: one using a beta-spectrometer in the classical ”target - detector” configuration, and the other relies an originally developed total absorption -spectrometer consisting of a pair of Si(Li)-detectors. These measurements correspond to the application of two approaches to determining the detector response function - accurate M-C modeling and the use of a detector response function close to a Gaussian function.
II.1 Spectrometer “target-detector” type
The “target-detector” type spectrometer was created on the basis of a semiconductor Si(Li) detector with a thickness of mm and the sensitive area diameter of mm, which had the “top-hat” geometry (Fig. 1(A)). These dimensions ensure an effective absorption of electrons with energies up to MeV.
The Si(Li)-detector produced from p-type single-crystal silicon with a resistivity of k and a carrier lifetime of s using standard technology that has been developed and tested by the Petersburg Nuclear Physics Institute. The characteristics of such detectors were described in [33, 34, 35].
The Si(Li)-detector was placed in a holder and equipped with a tungsten collimator with diameter of mm. The total insensitive layer of the input detector window including the gold-plated -contact was about 470 nm of silicon equivalent. The thickness of the insensitive layer of diffuse lithium on the back side of the detector was identified through the ratio of the total absorption peak areas for 241Am source gamma lines with energies of keV and keV to be .
The 144Ce source in the form of a dried drop of colloidal solution was placed on the surface of a maylar substrate with thickness of m. The substrate was fixed in a delrin ring and placed on a supporting structure at the distance of mm from the detector surface. The thickness distribution over the spot of the dried source 144Ce was determined from the measured energy spectrum of alpha particles originating from the decays of 241Am and 244Cm, which were preserved in the source as low activity impurities after purification of the 144Ce source. The average thickness of the source was determined as m.
The entire installation of the Si(Li)-detector and source was placed in a vacuum cryostat and was cooled down to the temperature of liquid nitrogen (LN).
The Si(Li)-detector was equipped with a charge-sensitive preamplifier with a field-effect transistor placed inside the vacuum cryostat and cooled down to the LN temperature. The negative bias voltage of kV was applied directly to the gold coating of the detector.
The signal from the preamplifier was registered by 8-channel CAEN v1725 digitizer with sampling rate of MHz. Further transformations with the signal were performed digitally. The time reference and the record trigger of the event were formed on the basis of the signal after digital CR-2RC formation, while the amplitude analysis was performed using triangular shaping. The signals were recorded on event-by-event basis, a time reference and amplitude were recorded for each event.
The setup also included a scintillation detector based on a 3” ( kg) bismuth orthogermanate (BGO) crystal coupled to a Hammamatsu R1307 PMT, the output of which was connected to the digitizer channel. Time and amplitude information was recorded for the signal of the BGO detector, which made it possible to analyze time-amplitude coincidences allowing to discriminate the spectrum of the allowed transition . The time resolution of the spectrometric paths was about ns.
The energy calibration was performed with help of total absorption peaks for gamma quanta with energies of keV and keV of the 207Bi source (Fig. 2). The energy resolution FWHM (full width at half maximum), determined by the conversion electron peak with an energy of keV, was keV ( eV).
Measurements were carried out continuously for 2024 hours, with short technological stops. The long duration of the measurements was defined by the low activity of the 144Ce source, which was Bk. The drift of the gain coefficient of the Si(Li) detector spectrometric channel during the long-term measurements was compensated by re-calibrations driven by the peak of conversion electrons with an energy of 91.5 keV at decays of 144Ce and the 5.902 MeV -peak of 244Cm, which were presented in the measured spectra.
The “target - detector” type spectrometer has a number of undoubted advantages, in particular, it allows one to measure -spectra of various sources using the same detector with quick replacement of the sources [35]. The main problem with the scheme is the backscattering of electrons from the crystal surface, which depends on the angle of incidence and leads to a large low-energy ”tail” in the detector response function.
II.2 4-spectrometer
The original spectrometer with 4 geometry was designed on the basis of two Si(Li) detectors with the thicknesses of about mm each, which exceeds the ionisation path of electrons with an energy of MeV. The detectors were specially manufactured for this experiment and they were made in the “top-hat” shape with external diameters of mm and mm, height of mm and diameters of the sensitive area of mm and mm. The Si(Li) detectors had different external diameters for the convenience of assembling the spectrometer with two detectors docked closely (Fig. 1, (B)) [36, 37, 38].
The characteristics of the detectors were tested in a separate vacuum cryostat using -, X-rays, conversion and Auger electrons from the source. The energy resolution measured for conversion electrons with energy of 481 keV from 207Bi source was keV for both detectors.
The total thickness of the insensitive layer of the Si(Li) detector, contributed by the deposited layers of palladium and gold and the surface layer of silicon, corresponds to a thickness of about 500 nm in silicon equivalent. Passing through such a thickness, electrons with energies of keV and MeV lose about keV and keV respectively.
Comparison of the measured intensities of X-ray Kα1, Kα2 and Kβ123 lead peaks with the results of Monte Carlo calculations using GEANT4.10.6 package allowed us to determine the thickness of the detector sensitive i-region. The useful thickness of both detectors determined in this way exceeds mm, which ensures, if one does not takes into account the backscattering, the complete absorption of any electrons with the energies less than MeV.
In the center of the front surface of one of the detectors a small cavity with a diameter of mm and a depth of mm was made by chemical etching. The 144Ce -source was placed into the cavity, directly onto the gold coating of the planar Si(Li) detector. The second detector was located on the detector with a cavity without any gap, and the bias voltage was applied to the resulting joined metal contact.
The assembly was placed in a vacuum cryostat and cooled down to the temperature of liquid nitrogen. The spectrometric channels were completely similar to the ones used in the “target-detector” type spectrometer and included preamplifiers with cooled field-effect transistors, the outputs of the preamplifiers were connected to the CAEN v1725 digitizer.
The total registered energy spectrum is obtained as the sum of the single event spectra from each of the Si(Li)-detectors and the total energy spectrum of events, recorded by both detectors in coincidence. The obtained spectrum solves the response function problem associated with backscattering of electrons from the detector surface.
The BGO-detector placed on the cryostat cap at a distance of mm from the source was also connected to one of the inputs of the digitizer.
The proposed full absorption original -spectrometer can be used for direct measurements of -specta and does not require corrections of the response function for the electrons backscattering from the crystal surface.
III DATA ANALYSIS
The nucleus undergoes three first-order forbidden non-unique beta-decays to the ground ( keV, %) and excited levels ( keV, % and keV, %) of the nucleus.
Decays of the to the ground ( keV, %) and first excited ( keV, %) states of the nucleus correspond to first-order forbidden non-unique and unique transitions, respectively. Decay into the level of with an energy of keV ( keV, %) is an allowed transition, the fitting of the beta spectrum does not require additional shape corrections as in the case of forbidden transitions [39].
The main objective of the work is to determine the spectrum of electron antineutrinos in the decay of with an energy above keV, the threshold of the inverse beta decay reaction on hydrogen. This is possible at electron energies less than keV in the case of a transition to the ground state and at an electron energy less than keV in the decay to the first excited state.
The energy spectra of particles arising from the decay of , which are -, Auger-, and conversion electrons, X- and -rays, have a complex form for fitting. Therefore, we will precisely fit the -spectrum of in the range of () keV, capturing a small region of the spectrum of near the end-point energy , in order to use it as an additional reference energy point. To determine the neutrino spectrum at neutrino energies greater than (), extrapolation of the obtained -spectrum to the low-energy region will be used.
In our work the detailed expression for electron energy spectrum of the given -transitions included the following components [40, 41, 42, 12, 43]:
where is a normalization constant, is electron momentum, is total electron energy in units of electron mass (, – kinetic energy of electron), corresponds to the endpoint energy of -spectrum, is electric charge (i.e. proton number) of the daughter nucleus, and is Fermi function.
The first part of expression (III) is phase space factor accounting for distribution of decay energy between produced particles (i.e. electron and antineutrino). The second factor is so called Fermi function describing the Coulomb interaction of outgoing electron with daughter nucleus. Fermi function originates from solution of Dirac equation in assumption of point-like and infinitely heavy nucleus and in fact represents the leading order QED correction for description of -decay.
In case of high precision measurements the basic Fermi approximation becomes insufficient and one has to take more intricate effects into account, which is achieved by introducing additional electromagnetic and weak corrections accounting for finite size and mass of the daughter nucleus () [40], screening of nuclear electric field by electrons of atomic shell () [42, 12], and radiative correction due to emission of photons by charged particles involved in -decay () [46, 43].
The screening correction accounts for the contribution of atomic electrons still remaining in “parent” configuration affecting the Coulomb potential of the “daughter” nucleus produced in -decay. The initial work around for this problem suggested simple rescaling of the total energy of a -particle [44], where is an effective shift of Coulomb potential due to the screening of atomic electrons. While this approach has been proven to be adequate for energies [45], due to the energy rescaling this technique becomes invalid in low-energy part of the spectrum when , exactly where it is of the most importance for evaluation of the high-energy end of corresponding neutrino spectrum.
In case of Ce and Pr atoms the screening shift MeV [42]. Since realistic description of screening correction below introduces significant complexity, we performed the upper estimate of potential uncertainty caused by screening effect by performing calculations for two border cases: no screening below () and maximum screening below (). These two options introduce the greatest uncertainty into the neutrino spectra and, as a consequence, into the expected count rate of the inverse beta decay events and were considered as the largest systematic uncertainty.
The radiative correction to the -spectrum at order has been computed in [46, 47, 48]. The radiative corrections have two components connected with the virtual photons on the Feynman lines of the charged particles and the emission of real internal bremsstrahlung (IB) photons [49]. As a result, when reconstructing the antineutrino spectrum from the -spectrum, the correction is replaced not by , but a new correction [49, 48]. The -spectrum and the antineutrino spectrum are described by two different radiative corrections and , respectively. This approach should be used when determining the antineutrino spectrum based on ”target-detector” spectrometer data since only % of the IB photons are emitted in the Si(Li)-detector solid angle.
The peculiarity of the described above -spectrometer is that it registers with high efficiency the IB photons with energy up to 100 keV, which are considered as part of the radiative corrections. For example, 1 MeV electron appearing in -decay emits 75% of the IB photons with an energy of less than 100 keV, of which only 6% leave the -detector without interaction. As a result, we used two different forms of the radiative correction function fitting the data of ”target-detector” and -spectrometer. For the -spectrometer we used the correction introduced for the neutrino spectrum [48, 49] with the natural replacement of by .
In addition to electromagnetic and kinematic effects, another higher order correction to Fermi function arise from internal nuclear structure and matrix elements of a given -transition. Nuclear shape-factor effectively accounts for complex collective nuclear interactions and becomes significant in case of forbidden transitions [43]. It is assumed that for allowed transitions , but except for some trivial cases of allowed and unique-forbidden transitions, analytical evaluation of nuclear shape-factor appears to be very difficult, thus leaving the experimental extraction the only viable option.
As noted above, we need to extrapolate the shape-factor to the region of electron energies below keV. The encouraging fact is that the expected theoretical dependence of for the first-order forbidden non-unique transition is almost constant in this region, smoothly increasing by only .
The weak magnetism correction is not included in equation (III) because it is vanished for the main decay branch of -nucleus [50].
III.1 Analysis of the data of spectrometer “target-detector” type
The measured spectrum is a convolution of the beta spectrum (III) and the function that describes the detector response to an electron with total energy , while is the energy recorded by the -spectrometer. The response of the detector was modeled by the Monte Carlo method using the GEANT4.10.6 package and G4Em_Standard_Physics_option4 package for the electromagnetic interactions, specially designed for detailed accounting of these interactions at low energies.
The simulation was carried out with a detailed description of the geometry of the installation and additional adjustment of the total thicknesses of the insensitive layers of the detector as well as the distribution of the thickness of the target, reproduced in the shape of the -spectra of and whose small impurities were present in the source. As an additional degree of freedom for back-scattered electrons, a variable ratio was introduced for the areas of the response function for electron energy below and above the original electron energy reduced by 25 keV ( keV).
The analysis of the “target - detector” experimental data was rather sophisticated as the duration of 144Ce source storage time by the moment of the experiment begin was already days, resulting thus in low activity and a complex composition of the background components in the measured spectrum. In order to solve this problem, the results of the long-term measurements were divided into two consecutive datasets. The analysis of the difference between the first and second parts of the dataset made it possible to significantly reduce the contribution of long-lived background components at the cost of losing statistics from the decays of 144Ce and 144Pr. As a result of the measurements, one-hour series were selected, which were divided into two equal parts. The spectrum corresponding to the first part was fitted by the sum of the second part spectrum, the beta spectra of 144Ce and 144Pr and the complementary background components (Fig. 3).
The main background component present in the measured spectrum was associated with the activity of the 154Eu (). The expected spectrum of 154Eu was simulated by the Monte Carlo method similarly to the response function. The spectrum measured over 1012 hours was fitted with the maximum likelihood method using the function (Fig. 3). The fitting was carried out in the range keV, which includes part of the -spectrum of . This made it possible to use the well-established -value ( keV) of 144Ce-decay.
The nuclear shape-factor is the function searched in the problem of the -spectrum measurement. The form factor was parameterized in powers of similarly to [51]:
where and are free parameters. As a result of the analysis, the following expression for the form factor function was obtained:
The accuracy of the obtained result (3) is limited by the weak activity of the 144Ce source, which requires a small distance from the detector to the target, that leads to large angles of incidence of electrons onto the crystal surface and an increased proportion of backscattered electrons. The accuracy in determination of the the nuclear form factor parameters and could be significantly increased by usage of a more active source with “target-detector” spectrometer.
The spectrum of the allowed transition , fitted similarly with the background component of random coincidences, has shown a good agreement of the model with the experiment (Fig. 4). The fit with fixed and gives the minimum -value is 288.6 for 256 degrees of freedom, which corresponds to the acceptable -value = 0.08. If the parameters and are free during the fit, their best-fit values are compatible with zero within their uncertainties.
III.2 Analysis of the data of 4-spectrometer
Neglecting the energy of the recoil nucleus, the energy of the electron and neutrino are related by a simple relation . The detector geometry makes it possible to measure the real spectrum of electrons and, ideally in the case of delta response , allows one to directly determine the antineutrino spectrum from the measured electron one without using the fitting procedure based on eq.(III). However, at electron energies where transitions to excited states of or decays of contribute to the measured spectrum, corrections must be made for the corresponding branching factors.
Although the response of the spectrometer is close to the Gaussian function, one should take into account the passage of the insensitive layers of the source and detectors. In order to account for these effects detailed M-C calculations were carried out for different electron energies.
The detector response function determined as a result of fitting looks like as a convolution of a Gaussian function with an exponential tail, which describes small energy losses in the target and non-sensitive layers of the detectors. As an example, for 1 MeV electrons the standard deviation for response function (the square root of the second central moment ) is 56.8 keV and the fraction of electrons detected with an energy less then 0.850 MeV is 3.4%.
The response function calculated by the M-C method was described analytically as:
The function was estimated from the condition of conservation of the full integral of the response function. describes the dependence of the standard deviation of the response function with two additional free parameters and : , where the dependence is obtained by the Monte Carlo method. The parameters , are responsible for imperfections in the modeling, is the Heaviside step function.
The spectrum of source measured by -spectrometer over 106 hours is shown in Fig. 5. The fitting was carried out by the maximum likelihood method with the likelihood function with the shape factor similar to the one (2) used in the data analysis of the “target-detector” spectrometer.
As a result of the analysis, the following values of the parameters and of the nuclear form factor function were obtained:
From (3) and (5) one could see that the results for the parameters and obtained with two different beta spectrometers agree with each other within the uncertainties (Fig.3 and Fig.5). The achieved accuracy of and determination in the measurement with the spectrometer is significantly higher than for the ”target detector” setup. The fitting of the allowed transition measured spectrum has shown a good agreement of the theoretical model used with the experiment as in the case of the measurement with the target-detector spectrometer (Fig. 6).
The -spectrum has been previously measured in several works [51, 52, 53, 54, 55] that used different expressions for the Fermi function and for the form factor. Therefore, a direct comparison of the results obtained is difficult. Nevertheless, our result for is more precise but consistent with [55, 54] within experimental uncertainties.
IV NEUTRINO SPECTRUM
Since the -spectrum shape measured by the -spectrometer turns out to be more accurate, it was used to determine the neutrino spectrum. Provided there is a measured shape of each component of -spectrum one could evaluate a neutrino spectrum though conservation of energy in the decay as the total energy of each decay is given by its endpoint energy. The neutrino spectrum is obtained from (III) by the replacements with .
The main interest of sterile neutrino experiments lies above the energy threshold of inverse beta decay on hydrogen, both in terms of shape and spectrum fraction. As for – source, only two -transitions contribute to the neutrino spectrum above this threshold: and with the branching ratios of % and %, respectively.
Fig. 7 illustrates the spectrum of electron antineutrinos from 144Pr source. The numerical values for neutrinos with energies above the IBD threshold are shown in the Tab. 1. The uncertainties of are calculated with toy Monte-Carlo method based on the uncertainties of the evaluated form-factor parameters and from (5) and taking into account their correlation. The parameters evaluated with this method were simulated 10000 times and used for performing 2-dimensional spectrum density plot. The RMS of density distributions of its monoenergetic projections was taken as the statistical uncertainty of the spectrum.
The uncertainty of the energy scale is determined by the accuracy of the -value of -transition of the to the ground state of [39].
Neutrinos from the decay of have the same intensity as -neutrinos, but their energies are less than 318 keV (the region to the left of line 1 in Fig. 7). The threshold value of the IBD reaction on hydrogen in the laboratory system is 1.806 MeV (line 2).
| E, keV | N, 1/MeV | E, keV | N, 1/MeV | E, keV | N, 1/MeV |
| 1809 | 0.54230(80) | 2209 | 0.49961(79) | 2609 | 0.35291(74) |
| 1819 | 0.54259(80) | 2219 | 0.49713(79) | 2619 | 0.34837(73) |
| 1829 | 0.54281(80) | 2229 | 0.49459(79) | 2629 | 0.34378(73) |
| 1839 | 0.54297(80) | 2239 | 0.49198(79) | 2639 | 0.33916(73) |
| 1849 | 0.54306(80) | 2249 | 0.48931(79) | 2649 | 0.33451(73) |
| 1859 | 0.54308(80) | 2259 | 0.48658(79) | 2659 | 0.32983(72) |
| 1869 | 0.54302(80) | 2269 | 0.48378(78) | 2669 | 0.32511(72) |
| 1879 | 0.54290(80) | 2279 | 0.48093(78) | 2679 | 0.32037(72) |
| 1889 | 0.54271(80) | 2289 | 0.47803(78) | 2689 | 0.31559(72) |
| 1899 | 0.54245(80) | 2299 | 0.47388(78) | 2699 | 0.31079(71) |
| 1909 | 0.54211(80) | 2309 | 0.46789(78) | 2709 | 0.30597(71) |
| 1919 | 0.54171(80) | 2319 | 0.46485(78) | 2719 | 0.30112(71) |
| 1929 | 0.54124(80) | 2329 | 0.46174(78) | 2729 | 0.29625(71) |
| 1939 | 0.54070(80) | 2339 | 0.45858(78) | 2739 | 0.29137(70) |
| 1949 | 0.54008(80) | 2349 | 0.45535(78) | 2749 | 0.28646(70) |
| 1959 | 0.53940(80) | 2359 | 0.45207(78) | 2759 | 0.28154(70) |
| 1969 | 0.53864(80) | 2369 | 0.44873(77) | 2769 | 0.27661(69) |
| 1979 | 0.53782(80) | 2379 | 0.44532(77) | 2779 | 0.27166(69) |
| 1989 | 0.53693(80) | 2389 | 0.44187(77) | 2789 | 0.26671(69) |
| 1999 | 0.53596(80) | 2399 | 0.43835(77) | 2799 | 0.26174(68) |
| 2009 | 0.53492(80) | 2409 | 0.43479(77) | 2809 | 0.25677(68) |
| 2019 | 0.53382(80) | 2419 | 0.43116(77) | 2819 | 0.25180(68) |
| 2029 | 0.53264(80) | 2429 | 0.42749(77) | 2829 | 0.24683(67) |
| 2039 | 0.53139(80) | 2439 | 0.42375(76) | 2839 | 0.24186(67) |
| 2049 | 0.53008(80) | 2449 | 0.41997(76) | 2849 | 0.23691(67) |
| 2059 | 0.52869(80) | 2459 | 0.41614(76) | 2859 | 0.23196(66) |
| 2069 | 0.52723(80) | 2469 | 0.41225(76) | 2869 | 0.22703(66) |
| 2079 | 0.52571(80) | 2479 | 0.40831(76) | 2879 | 0.22212(65) |
| 2089 | 0.52411(80) | 2489 | 0.40432(76) | 2889 | 0.21723(65) |
| 2099 | 0.52244(80) | 2499 | 0.40028(76) | 2899 | 0.21238(65) |
| 2109 | 0.52071(80) | 2509 | 0.39620(75) | 2909 | 0.20757(64) |
| 2119 | 0.51891(79) | 2519 | 0.39207(75) | 2919 | 0.20280(64) |
| 2129 | 0.51703(79) | 2529 | 0.38789(75) | 2929 | 0.19809(64) |
| 2139 | 0.51509(79) | 2539 | 0.38367(75) | 2939 | 0.19345(63) |
| 2149 | 0.51308(79) | 2549 | 0.37940(75) | 2949 | 0.18888(63) |
| 2159 | 0.51100(79) | 2559 | 0.37509(75) | 2959 | 0.18442(62) |
| 2169 | 0.50886(79) | 2569 | 0.37073(74) | 2969 | 0.18011(62) |
| 2179 | 0.50665(79) | 2579 | 0.36634(74) | 2979 | 0.17610(62) |
| 2189 | 0.50437(79) | 2589 | 0.36190(74) | 2989 | 0.17327(62) |
| 2199 | 0.50202(79) | 2599 | 0.35743(74) | 2995 | 0.22220(80) |
Considering the precision given by our measurement one could apply toy Monte-Carlo to convert the uncertainties on the form-factor parameters to the fraction of the -spectrum above the threshold IBD reaction that is so the statistic precision of this value is as low as 0.01 %. The systematic uncertainty evaluated as the difference of the values derived for two boundary cases of low-energy screening correction is 0.00022 (0.044 %). Greater systematic error is associated with the uncertainty of the energy scale at 1.806 MeV that lead to 0.0065 (0.13 %).
The expected number of IBD events is determined by the integral of the product of the reaction cross section and the neutrino flux (Fig.8). Using the cross section of IBD reaction [56, 57, 58, 59], one could evaluate the experimental neutrino count rate, that is cm2 per one decay of nucleus. The errors reported here are due only to uncertainties in the neutrino spectrum of connected with the shape (mainly due to the uncertainty in the screening correction 0.0041 (0.09%) and energy scale . The latter can be substantially improved by any refinement of the -value of -decay. The obtained precision of in % would correspond to an experimental uncertainty of neutrino count rate expected for a sterile neutrino experiment using Pr source. In this case this value would be sub-dominated by other experimental uncertainties, e.g. the precision of the source calorimetry for Borexino-SOX experiment has reached precision of % [30].
We can also provide an analytical interpolation of the spectrum, that could be expressed in two energy intervals. For neutrino energies above the unique forbidden transition endpoint of MeV, the spectrum is interpolated as follows:
Between the inverse beta decay threshold of MeV and MeV, we also include a transition to the excited level of in 144Nd, adding an additional component :
where the parameters are defined according to the Tab. 2. The accuracy of this analytical interpolation could be estimated by the ratio of two integrals: the product of the measured spectrum with the cross section of the inverse beta decay and the product of the interpolation function with the same cross section. For neutrino energies greater than MeV, the deviation of these two values was found to be as low as %.
V CONCLUSION
The -spectra of the source were studied using two spectrometers based on semiconductor Si(Li) detectors. The spectra were measured using a beta spectrometer in the classical “target-detector” scheme and with an original full absorption spectrometer consisting of two of Si(Li) detectors. The models of the beta spectrometer response function calculated through the Monte Carlo simulation were tested by fitting the spectrum of the allowed beta transition and have shown good agreement between the response models and experimental data. The function of the nuclear form factor of the ground state beta transition in has been defined as . The electron antineutrino spectrum of the decays was obtained from the measured -spectrum and the form factor parameters. The reduced cross section for the inverse beta decay reaction for the electron anti-neutrino source is cm2 per which is the most accurate measurement up to date.
ACKNOWLEDGEMENTS
This work has been supported by Russian Science Foundation (project no.24-12-00046).